Advantage Tennis Coaching in Brisbane offers coaching for junior, elite and mature tennis players. Nicholas Cash, the sole proprietor, has identified that the business cycle will be seasonal, with most activities in the spring and summer months. As a result, Nicholas is predicting net working capital will be negative from the beginning of April to the end of September. Nicholas arranges with his bank for funds to finance net working capital during this period. He requires 180-day money and about $81,000. His bank offers commercial bills at 5 per cent yield, including fees.

How much does Advantage Tennis Coaching receive? (Round answer to 0 decimal places, e.g. 5,275.)
Answer #1

Working Capital loan amount = $81,000

Rate = 5%

Tenure = 180 days

Interest charged by Bank = (Principal X Rate X Tenure / 365)

= (81000 X 5% X 180/365) = (4050 X 180/365) = 1997

Hence, the amount received on availing the working capital loan = Principal amount - Interest paid = 81000 - 1997 = $79,003